OpenStudy (anonymous):

Solve by factoring. 4x^ – 4x – 3 = 0

OpenStudy (anonymous):

4x^2 - 6x +2x - 3 = 0 2x(2x - 3) +1 (2x - 3) = 0 (2x-3)(2x+1)=0

OpenStudy (anonymous):

thank you :)